OpenStudy (anonymous):

Find derivative of y=tan^-1sqrt(x^2-1)+csc^-1x, x>1Find derivative of y=tan^-1sqrt(x^2-1)+csc^-1x, x>1@Mathematics

OpenStudy (anonymous):

\[\frac{d}{dx}\left(\tan ^{-1}\left(\sqrt{x^2-1}\right)+\csc ^{-1}(x)\right)=\frac{1}{x \sqrt{x^2-1}}-\frac{1}{\sqrt{1-\frac{1}{x^2}} x^2} \]Go to WolfFramAlpha.com and type in the following: derivative of tan^-1 sqrt (x^2 - 1) + csc^-1 x then click on "Show steps"